OpenStudy (anonymous):

surface area of revolved surface between x=sqrt(x^2+1) x=0 x=3 about the y-axis

OpenStudy (anonymous):

im sorry its the x-axis not the y-axis

OpenStudy (freckles):

have you found dx/dy yet?

OpenStudy (freckles):

wait

OpenStudy (freckles):

I'm confused... I though you had x=sqrt(y^2+1)

OpenStudy (freckles):

but you have x=sqrt(x^2+1)

OpenStudy (freckles):

this is an equation not a relation between x and y

OpenStudy (anonymous):

oh sorry its y= sqrt( x^2 +1) typo on my part

OpenStudy (freckles):

ok so have you tried to find dy/dx?

OpenStudy (anonymous):

i got x/sqrt(x^2+1)

OpenStudy (freckles):

and also if it is about the x-axis then we have: \[\int\limits 2 \pi y ds \\ \text{ where you can make either choice in } ds \text{ :} \\ ds=\sqrt{1+(y')^2} dx \text{ if } y=f(x), a \le x \le b \\ \text{ or } \\ ds=\sqrt{1+(x')^2} dy \text{ if } x=h(y), c \le y \le d\] checking your y'... \[y=\sqrt{x^2+1} \\ y^2=x^2+1 \\ 2y dy=2x dx \\ y dy =x dx \\ y \frac{dy}{dx}=x \\ \frac{dy}{dx}=\frac{x}{y} \\ \frac{dy}{dx}=\frac{x}{\sqrt{x^2+1}} \] that is right

OpenStudy (freckles):

\[\int\limits_0^3 2 \pi y \sqrt{1+(\frac{x}{\sqrt{x^2+1}})^2} dx \\ \\ \int\limits_0^3 2 \pi \sqrt{x^2+1} \sqrt{1+(\frac{x}{\sqrt{x^2+1}})^2} dx \]

OpenStudy (freckles):

are you able to complete the problem or are you getting stuck somewhere

OpenStudy (anonymous):

im lagging bad

OpenStudy (anonymous):

yeah im stuck on what to do next

OpenStudy (freckles):

can you show me what you have so far?

OpenStudy (freckles):

or were you not able to do anything from what I last wrote?

OpenStudy (anonymous):

yeah i couldn't get past that last part

OpenStudy (freckles):

\[\int\limits_0^3 2\pi \sqrt{x^2+1} \sqrt{1+\frac{x^2}{x^2+1}} dx \\ \int\limits_0^3 2\pi \sqrt{x^2+1} \sqrt{\frac{x^2+1}{x^2+1}+\frac{x^2}{x^2+1}} dx \\ \int\limits_0^3 2 \pi \sqrt{x^2+1} \sqrt{\frac{x^2+1+x^2}{x^2+1}} dx \\ \int\limits_0^3 2 \pi \sqrt{x^2+1 } \sqrt{\frac{2x^2+1}{x^2+1} } dx \\ \int\limits _0^3 2 \pi \frac{\sqrt{x^2+1}}{\sqrt{x^2+1}} \sqrt{2x^2+1} dx \\ \int\limits_0^3 2 \pi \sqrt{2x^2+1} dx\] what about now?

OpenStudy (anonymous):

i plugged it into wolfram and got this constant+pi sqrt(2 x^2+1) x+(pi sinh^(-1)(sqrt(2) x))/sqrt(2)

OpenStudy (anonymous):

but it doesnt look right to me

OpenStudy (freckles):

try a trig substiution recall \[\tan^2(\theta)+1=\sec^2(\theta)\]

OpenStudy (anonymous):

ok, srry but i have to go, i have class

OpenStudy (freckles):

\[2x^2+1 =(\sqrt{2}x)^2+1 \\ \text{ comparing } \\ (\sqrt{2} x)^2+1 \text{ with } (\tan(\theta))^2+1 \\ \text{ you should want to make the substitution } \sqrt{2} x=\tan(\theta)\]